OpenStudy (anonymous):

Katrina is fencing her yard, which is shaped as a rhombus. She used (5x + 1) meters of fencing material for one side and (7x - 11) meters for the adjacent side. What is the perimeter of the yard? Answer 116 meters 124 meters 140 meters 168 meters

OpenStudy (anonymous):

7x - 11 = 5x + 1 7x -5x = 1 + 11 2x = 12 x = 6 2(7(6)) - 11) + 2(5(6)) + 1) = Perimeter 62 + 62 = 124 units

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

yes it does alot thank u for showing me how to do it

OpenStudy (anonymous):

Your welcome :D

OpenStudy (anonymous):

*You're ;D